OpenStudy (anonymous):

Find all the intervals on which f(x) = x^2 + 1 / x^2 is concave upward

OpenStudy (anonymous):

1) find the second derivative i will wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6x^-4

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

the first derivative is -2x^-3

OpenStudy (anonymous):

what happened to the \(x^2\) part? did it disparu?

OpenStudy (anonymous):

you were close, you just forgot the \(2\)

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

-2 times -3 is 6

OpenStudy (anonymous):

\[f(x)=\color{red}{x^2}+\frac{1}{x^2}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you seem to be ignoring the poor \(\color{red}{x^2}\) part what did it do to you?

OpenStudy (anonymous):

I am not sure what you mean

OpenStudy (anonymous):

i used the quotient rule

OpenStudy (anonymous):

you gotta take the second derivative of that as well

OpenStudy (solomonzelman):

you take the second derivative of each term.

OpenStudy (anonymous):

yes using the quotient rule I got f'(x) = -2x^-3 and f''(x) = 6x^-4

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (anonymous):

\[y=x^2\\ y'=2x\\ y''=2\]

OpenStudy (anonymous):

you got the second derivative right for the \[\frac{1}{x^2}\] part, but you forgot about the \(x^2\) part

OpenStudy (anonymous):

2x-2x^-3 is the first derivative and the secon is 2+6x^-4

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

see what i meant about the 2?

OpenStudy (anonymous):

yeah haha

OpenStudy (anonymous):

now it is nice to use exponential notation to find derivatives, but to compute you need positive exponents \[f''(x)=2+\frac{6}{x^4}\]

OpenStudy (anonymous):

and your job is to find out where this beast is positive, and where it is negative, i.e. for which values of x is it positive or negative, which we do with our eyeballs

OpenStudy (anonymous):

how about \(x^4\) is that ever negative?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and therefore neither is \(\frac{6}{x^4}\)

OpenStudy (anonymous):

and if \(\frac{6}{x^4}\) is never negative, if you add 2 it will not be negative either (remember that 2?)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so the second derivative is always positive, (unless it its undefined)

OpenStudy (anonymous):

making your original function \[f(x)=x^2+\frac{1}{x^2}\] always concave _________

OpenStudy (triciaal):

which is the original